OpenStudy (anonymous):

prove sin^2x+cos^2x=1

OpenStudy (lalaly):

mmmm.....

OpenStudy (amistre64):

i thought pythag did that already

OpenStudy (anonymous):

help

OpenStudy (anonymous):

lol ask this to mimi x3

OpenStudy (anonymous):

yes mimi boosted 3 times will help me if u dont help me amistre

OpenStudy (anonymous):

it come from the Pythagor theorem

OpenStudy (anonymous):

but how ???????????

OpenStudy (amistre64):

x^2 + y^2 = r^2\[\frac{x^2}{r^2}+\frac{y^2}{r^2}=1\] \[(\frac{x}{r})^2+(\frac{y}{r})^2=1\] sub in the defs for sin and cos

OpenStudy (anonymous):

thnx amistre

OpenStudy (amistre64):

youre welcome :)